suppose you start saving today for a $20000 down payment that you plan to make on a house in 10 years. Asssume that you make no deposits into the account after your intial deposit. For each account described, how much would you have to deposit now to reach your $20000 goal in 10 years? An account with daily compounding and an APR of 4%
Answer by rfer(12730)

P=A/(1+.04/365)^365*10
P=20000/(1.000109589)^3650
P=20000/1.49179178
P=$13406.70